# Generalized Hypercube Queuing Models with Overlapping Service Regions

### v2

@@
Abstract
We present a generalized hypercube queue model, building upon the original model by Larson, focusing on its application to overlapping service regions such as police beats. To design a service region, we need to capture the workload and police car operation, a type of mobile server. The traditional hypercube queue model is constrained to light-traffic operations, where officers are either “busy” or “not busy”. However, contemporary service operations often experience saturation, necessitating the inclusion of heavy-traffic operations with queue lengths greater than one. The motivation to address the heavy-traffic regime stems from the increased workload and staff shortages prevalent in modern service systems. The design of overlapping regions is inspired by boundary effects in crime incidents, which require overlapping patrol regions for efficient resource allocation. Our proposed model addresses these issues using a Markov model with a large state space represented by integer-valued vectors. By leveraging the sparsity structure of the transition matrix, where transitions occur between states whose vectors differ by 1 in the $\ell_{1}$ distance, we can accurately solve the steady-state distribution of states. This solution can then be used to evaluate general performance metrics for the service system. We demonstrate promising numerical results of our model by simulating various types of artificial service systems. We also analyze the overlapping districting plans for the Atlanta police system with real 911 calls-for-service data using our model and discuss their implications for the police operation. Although in the paper we focus on police districting applications, the generalized hypercube queue model is applicable for other mobile server modeling for the general setup.
Keywords: Hypercube queuing model, Overlapping service regions
@@ 1 Introduction
Service systems, such as Emergency Service Systems (ESS), are vital in providing emergency aid to communities, protecting public health, and ensuring safety. However, these systems often face resource limitations, including constrained budgets and personnel shortages. As a result, the public and associated organizations need to evaluate and improve the efficiency of these systems to maintain their effectiveness.
@@
Service systems are generally designed such that a specific emergency response unit is primarily responsible for a small geographical area while remaining available to aid neighboring regions if necessary, either through a structured dispatch hierarchy or on an ad-hoc basis. For instance, in policing service systems, personnel are typically assigned to a designated area known as a beat. Police departments often allocate patrol forces by dividing a city’s geographical areas into multiple police zones. The design of these patrol zones significantly impacts response times to 911 calls. Officers patrol their assigned beats and respond to emergency calls, with each spatial region having one queue.
@@
A key component for analyzing service systems is to model the queueing dynamic of mobile servers, such as police cars, ambulances, and shared rides, depending on the context. Mobile servers refer to resources or servers that are not fixed in one location but are mobile, often to provide services in varying locations or under changing conditions. Some examples of mobile servers include ambulances, fire rescue trucks, on-demand delivery systems, and disaster response units. In his seminal work, Larson [13] introduced a general model for mobile servers, the so-called hypercube queue models. The model incorporated geographical components, or atoms, each with unique arrival rates, and mobile servers with distinct service rates. This model can capture the light traffic regime of police operations, compute the steady-state distribution of binary states efficiently, and conveniently evaluate general performance metrics.
@@
The motivation for revising and generalizing the classic hypercube model for police service systems stems from various factors, such as modeling police operations in heavy workload regimes (correlated with staff shortages), severe staff shortages, and the observed boundary effect of crime incidents. These observations are supported by data, as illustrated in Figures 1, 2, and 3. Each factor will be discussed in greater detail below.
@@
Figure 1: The increasing average police response time in Atlanta by month, from 2013 onwards.
@@
Figure 2: Operations statistics calculated using 911 data from the Atlanta Police Department, from 2013 to 2022, averaged for each of the 80 beats: distribution of hourly incident rate per beat, processing data in minutes, travel time in minutes, workload (the average time an officer need to travel to and process a single incident).
@@
Figure 3: The border effect: Burglary incidents reported by the Atlanta police department in 2017 (represented by red dots) clustered near the border between beats (indicated by black lines).
First, “busy” police operations critically require a flexible queuing model capable of capturing a regime not considered by the hypercube queue model. Despite its elegance, the hypercube queue model is limited in that it only captures the “light-traffic” regime, wherein officers in the beat are either “busy” or “not-busy”, and assumes that there is rarely more than one incident waiting in each server’s queue. However, in modern times, police operations systems are often in a saturated state, which differs from the “light-traffic” regime in the traditional hypercube queue model. Therefore, to understand the system in “heavy-traffic” where the queue length is larger than one, we aim to capture the distribution of states for each server’s queue length, rather than simply “busy” and “non-busy” states (represented by binary vectors) as in Larson’s classic hypercube queue models.
@@
Second, the traditional “one-officer-per-beat” operational mode has become increasingly difficult to implement due to severe staff shortages, which currently pose a significant challenge for many police departments, including those in the Atlanta metro area [7]. For example, as of January 2023, suburban Atlanta’s Gwinnett County (GA) reportedly “employed 690 officers out of an authorized strength of 939, leaving more than a quarter of sworn officer positions vacant” [22]. This leads to busier police operations and challenges in placing at least one officer per beat each day. To address officer shortages in practice, some municipalities have implemented flexible service regions in the middle of their districts, enabling officers to respond to incidents more effectively. This empirical strategy has proven successful. Moreover, by utilizing restricted service regions with some overlap, departments can balance full flexibility and complete isolation, ensuring officers can assist neighboring areas when needed without compromising the efficiency of their primary service areas.
@@
Third, implementing flexible service regions with overlapping beats can help address the so-called “border effect.” For example, in Atlanta, we have observed a notable boundary effect in the distribution of police incidents, with crime rates tending to be higher at the peripheries of police beats (refer to Figure 3 for a burglary case example). Additionally, response times at beat boundaries are often longer due to increased travel distances. Police officers also report that more criminal activity occurs near the borders of two beats, resulting in more 911 calls in these areas. The boundary effect may contribute to increased crime rates near the edges of designated police zones, underscoring the need for a more efficient allocation of law enforcement resources. This observation suggests that the proposed overlapping beat structure may offer the demonstrated theoretical advantages and potential additional benefits in combating the border effect.
Due to the aforementioned motivations, we present a generalized hypercube queue model in this paper. Our contributions are three-fold: (1) We generalize the classic hypercube queue model by Larson [13] to capture both the light traffic and heavy traffic, enabling us to characterize saturated systems with queue lengths greater than one explicitly; (2) We develop an efficient computational method to estimate the steady-state distribution for systems with a large number of servers; (3) We demonstrate the application of this model to represent the so-called overlapping patrol, where multiple police officers may respond to incidents at the boundaries of patrol regions.
(a) Hypercube
(b) Hyperlattice
Figure 4: Illustration of the differences in state spaces: (a) the traditional Hypercube model and (b) our proposed Hyperlattice model.
@@
To achieve our goal, we consider a Markov model with a potentially large state space represented by integer-valued vectors (in contrast, the hypercube queue model employs binary-valued state vectors, as shown in Figure 4). The sparse transition matrix allows for transitions between states whose vectors differ by 1 in the $\ell_{1}$ distance (in contrast, the hypercube queue model considers Hamming distance one transition for binary-valued state vectors). Based on this, we develop an efficient state aggregation technique to relate the generalized hypercube queue model to the known properties of a general birth-and-death process with state-dependent rates. This reduces the detailed balancing equations to a much smaller linear system of equations, enabling efficient computation for steady-state state distributions, which are key in evaluating general performance metrics as recognized in Larson’s original work [13]. Finally, we validate the effectiveness of the proposed framework by testing it with simulated cases and real-world 911 service call data from Atlanta, Georgia.
@@
This paper is organized as follows: We begin with a review of relevant literature on spatial queuing systems and Markov models with similar structures. In Section 2, we introduce a novel queuing framework for evaluating districting plans with overlapping regions. Section 2.1 discusses the problem setup and key assumptions. Section 2.2 presents a generalized hypercube queuing model – the hyperlattice model – for capturing system dynamics, while Section 2.3 introduces an efficient method for estimating its steady-state distribution. Section 2.4 introduces two key performance measures using the hyperlattice model. Lastly, we demonstrate the effectiveness of our framework through experiments on synthetic systems in Section 3 and conclude in Section 4, discussing potential extensions and generalizations.
1.1 Related works
@@
Spatial queuing systems modeling with overlapping service regions have received limited attention in the literature. However, recent developments in service systems and their growing complexity have spurred an increased interest in this field, particularly in the quest for more efficient dispatch policies. For instance, the concept of overlapping zones is discussed in [4], which only considers completely overlapping service regions. This approach may need to address higher crime rates near boundaries. In contrast, our study focuses on partially overlapping service regions, which are more challenging to analyze but better suited for our purposes.
@@
Richard Larson’s paper [13] and his co-authored book with Asaf Odoni, “Urban Operations Research” [14], introduced the hypercube queuing model, a pioneering work in this field. The hypercube queuing model offers a spatial queuing framework for analyzing and evaluating service systems’ performance. It permits servers to cross over into each other’s territories and provide aid when busy, making it particularly useful for emergency service systems like police and ambulance operations [2, 9, 19, 24, 25, 15]. However, it assumes identical service from all servers sharing a single queue, which may be oversimplified for real-world service systems and unable to capture more complex systems. In our context, servers may have limited authorized movement, resulting in separate queues for each server. This is because other servers cannot travel to specific regions to offer assistance. Consequently, a more flexible model may be necessary to represent such service systems accurately.
@@
Among existing work extending the classic hypercube queuing model for specific applications, a notable contribution is [5], which introduces a variation of the hypercube model. Their extension is designed to represent a dispatch policy in which advanced equipped servers serve solely life-threatening calls (called dedicated servers). Our model, however, differs as it allows servers to operate within the territories of other servers, presenting a distinct problem framework. In addition, they grouped servers with similar characteristics in their state space representation, aiming to lower computational demands. In our method, each server is still represented individually in the state space, managing to do so without significant computational expense. We also assume each server has its own queue, diverging from most Hypercube model variations that use a single queue for all servers.
@@
Recently, [20] introduced a flexible queue architecture for a multi-server model, comprising $n$ flexible servers and $n$ queues connected through a bipartite graph, with each server having its own associated queue. The paper mainly focuses on the model’s theoretical aspects and addresses the essential benefits of flexibility and resource pooling in such queuing networks. They specifically concentrate on the scaling regime where the system size $n$ tends to be infinite while overall traffic intensity remains constant. However, their model cannot be directly extended to our scenario, as it does not consider the geographical relationship among servers.
Other related works analyze service systems using similar modeling techniques [8]. For example, [3] proposes a new approximate hypercube spatial queueing model similar to the hyperlattice model, allowing multiple servers to be located at the same station and multiple servers to be dispatched to a single call. The state space of their model can be represented by a vector of non-negative integer, where the $i$-th integer indicates the number of calls in the system that are being served by the $i$-th server. Rather than reducing to a birth-death process, they use the M[G]/M/s/s queueing model to derive approximate formulas for the hypercube spatial queueing outputs. In [11], the paper describes a model developed to represent patrol car operations accurately. It is a multi-priority queuing model explicitly reflecting multiple car dispatches. Instead of focusing on overlapping patrol, this paper aims to provide a better basis for efficient patrol car allocation and enables police officials to evaluate policies, such as one-officer patrol cars, that affect the number of cars dispatched to various incident types. In [21], the authors examine similar systems where each server can handle a specific subset of “job types”, which can be seen as calls received in different regions. The proposed solution retains the identity and position of busy servers but does not indicate the particular job type being handled. Specifically, it only tracks the number of jobs between busy servers without specifying the sequence of job types. However, like the Hypercube queuing model, this model has a single queue where incoming jobs without feasibly available servers wait and are processed in a First-Come-First-Served order. As a result, this model is not directly applicable to represent service systems of interest with multiple queues in overlapping service regions. A recent paper [23] studies a distinct server routing-scheduling problem in a distributed queueing system, unlike the problem we are examining in this paper. This system consists of multiple queues at different locations. In such a system, servers are allocated across multiple queues and move among these queues to meet demands that are both random and fluctuate over time. Another paper [1] proposes a similar approach, modeling the system as a Markov process but delaying the discovery of a job’s type until a server needs to know the type. In comparison, our system does not differentiate calls (or job types) in any server. Another key difference is that this paper aims to derive exact asymptotics of stationary distributions from analyzing production contexts where the fluid limit of excursions to the increasingly rare event is nonlinear.

### v4

@@
[…]
Generalized Hypercube Queueing Models with Overlapping Service Regions
[…]
We present a generalized hypercube queueing model, building upon the original model by Larson (1974), focusing on its application to overlapping service regions, such as police beats. To design a service region, we need to capture the workload and police car operation, a type of mobile server. The traditional hypercube queueing model excels in capturing the dynamics of systems with light traffic, as it primarily considers whether each server is busy or idle. However, contemporary service operations often experience saturation, in which each server in the system can only process a subset of calls, and a queue in front of each server is allowed. Hence, the simple binary status for each server becomes inadequate, prompting the need for a more intricate state space analysis. Our proposed model addresses this challenge by utilizing a Markov model with a large state space, represented by non-negative integer-valued vectors, along with a truncated hyperlattice queueing model approximation. By exploiting the sparsity of the transition matrix, we efficiently compute the steady-state distribution of the truncated hyperlattice model and show that it provides a close approximation under canonical dispatching policies. This solution can then be used to evaluate general performance metrics for the service system. We validate the effectiveness of our model through simulations of various artificial service systems. We also apply our model to the Atlanta police operational system, which faces challenges such as an increased workload, significant staff shortages, and the impact of boundary effects on crime incidents. Using real 911 calls-for-service data, our analysis indicates that a police operations system with permitted overlapping patrols can significantly mitigate these problems, leading to more effective deployment of the police force. Although the paper focuses on police districting applications, the generalized hypercube queueing model is applicable to other mobile server models in the general setup.
[…]
Hypercube queueing model; Overlapping service regions
@@ 1 Introduction
Service systems, such as Emergency Service Systems (ESS), are vital in providing emergency aid to communities, protecting public health, and ensuring safety. However, these systems often face resource limitations, including constrained budgets and personnel shortages (Peng et al. 2020, Yoon et al. 2021, Zhu et al. 2022). As a result, the public and associated organizations need to evaluate and improve the efficiency of these systems to maintain their effectiveness.
@@
Service systems are generally designed such that a specific emergency response unit is primarily responsible for a small geographical area while remaining available to aid neighboring regions if necessary, either through a structured dispatch hierarchy or on an ad-hoc basis. For instance, in policing service systems, personnel are typically assigned to a designated area known as a beat. Police departments often allocate patrol forces by dividing a city’s geographical areas into multiple police zones. The design of these patrol zones has a significant impact on response times to 911 calls. Officers patrol their assigned beats and respond to emergency calls, with each spatial region having one queue.
@@
A key component for analyzing service systems is to model the queueing dynamics of mobile servers, such as police cars, ambulances, and shared rides, depending on the context. Mobile servers refer to resources or servers that are not fixed in one location but are mobile, often to provide services in varying locations or under changing conditions. Some examples of mobile servers include ambulances, fire rescue trucks, on-demand delivery systems, and disaster response units. In his seminal work, Larson (1974) introduced a general model for mobile servers, known as the hypercube queueing models. The model incorporated geographical components, or atoms, each with unique arrival rates, and mobile servers with distinct service rates. This model can capture the light traffic regime of police operations, compute the steady-state distribution of binary states, and conveniently evaluate general performance metrics.
@@
The motivation for this study stems from various factors, including modeling police operations with heavy workload (often linked to staff shortages), severe staff shortages, and the observed boundary effect of crime incidents. These observations are supported by data, as illustrated in Figures 1, 1, and 1. Each factor will be discussed in greater detail below.
[…]
@@
The increasing average police response time in Atlanta by month, from 2013 onwards.
[…]
@@
Operations statistics calculated using 911 data from the Atlanta Police Department, from 2013 to 2022, averaged for each of the 80 beats. These figures show that calls are coming in faster than they can be handled by the police, as indicated by the workload in minutes and the hourly rate of the calls.
[…]
@@
The border effect. Burglary incidents reported by the Atlanta police department in 2017 (represented by red dots) clustered near the border between beats (indicated by black lines).
First, modern police operations, often strained by high workload, necessitate a more flexible queueing model than what is provided by the traditional hypercube queueing model. The hypercube model, while elegant in its simplicity, is limited to representing each server’s status as simply “busy” or “idle” and operates under the assumption of a singular queue for the entire system. However, in today’s context, police operations systems frequently face overwhelming demands, leading to a scenario where all servers are constantly busy, which may not be accurately captured by the hypercube model. Therefore, to effectively understand systems with heavier loads, where there could be multiple queues and each queue might exceed a length of one, our objective is to model the distribution of queue lengths for each server. This approach goes beyond the basic “busy” and “idle” states (represented by binary vectors) used in Larson’s original hypercube queueing models, offering a more detailed and realistic representation of contemporary police operations.
@@
Second, the traditional “one-officer-per-beat” operational mode has become increasingly difficult to implement due to severe staff shortages, which currently pose a significant challenge for many police departments, including those in the Atlanta metro area (Cummings 2022). For example, as of January 2023, suburban Atlanta’s Gwinnett County (GA) reportedly “employed 690 officers out of an authorized strength of 939, leaving more than a quarter of sworn officer positions vacant” (Writer 2023). This leads to busier police operations and challenges in assigning at least one officer to each beat each day. To address officer shortages in practice, some municipalities have implemented flexible service regions within their districts, enabling officers to respond to incidents more effectively. This empirical strategy has proven successful. Moreover, by utilizing restricted service regions with some overlap, departments can strike a balance between full flexibility and complete isolation, ensuring that officers can assist neighboring areas when needed without compromising the efficiency of their primary service areas.
@@
Third, implementing flexible service regions with overlapping beats can help address the so-called “border effect.” For example, in Atlanta, we have observed a notable boundary effect in the distribution of police incidents, with crime rates tending to be higher at the peripheries of police beats (refer to Figure 1 for a burglary case example). Additionally, response times at beat boundaries are often longer due to increased travel distances. Police officers also report that more criminal activity occurs near the borders of two beats, resulting in a higher volume of 911 calls in these areas. The boundary effect may contribute to increased crime rates near the edges of designated police zones, underscoring the need for a more efficient allocation of law enforcement resources. This observation suggests that the proposed overlapping beat structure may offer the demonstrated theoretical advantages and potential additional benefits in combating the border effect.
Motivated by these operational realities, this paper introduces a new queueing framework that generalizes the classical hypercube model to accommodate overlapping service regions, individual queue lengths, and flexible dispatch policies. While overlapping patrol strategies have gained traction in practice, there is currently no mathematical framework that systematically evaluates their performance. Our proposed model addresses this gap and offers the following core contributions:
  1. 1.
    Hypercube queueing models for spatial service systems: We introduce a state-space representation that captures per-server queue lengths and structured service overlaps in spatial service systems, allowing analysis of system dynamics under light and heavy traffic.
  2. 2.
    Efficient steady-state computation for large-scale systems: We introduce a hyperlattice structure that, despite its infinite state space, enables tractable and efficient computation through structured sparsity and state truncation techniques.
  3. 3.
    Application to real-world patrol systems: We validate the model using data from Atlanta Police Department operations, demonstrating how overlapping patrol regions can alleviate operational strain and reduce the adverse effects of boundaries. Our framework offers actionable insights into how resource allocation and dispatch policies influence system performance.
[…]
Illustration of the differences in state spaces. Our model represents the status of each server based on the number of servers waiting in its corresponding queue. In contrast, the hypercube model reduces this to a single tail state when all servers are busy. As a result, our model’s state space forms a “hyperlattice”, expanding the original finite-state hypercube structure into an infinite space. While similar hyperlattice structures have been proposed in prior works, our approach utilizes them to represent a different state space.
@@
To achieve our goal, we consider a Markov model with a potentially large state space represented by integer-valued vectors, unlike the hypercube queueing model, which uses binary-valued state vectors as shown in Figure 1. In this model, each server has its own queue, and the integer in the state represents the number of calls waiting in that queue. Crucially, the policy that manages calls in each queue is determined by a hypergraph that reflects the overlapping nature of these regions. The design of the state space results in a sparse transition matrix that allows for transitions between states whose vectors differ by 1 in the $\ell_{1}$ distance (in contrast, the hypercube queueing model considers Hamming distance one transition for binary-valued state vectors). Based on this, we develop an efficient state truncation technique to relate the generalized hypercube queueing model to the known properties of a general birth-and-death process with state-dependent rates. This reduces the detailed balancing equations to a much smaller linear system of equations, enabling efficient computation for steady-state distributions, which are key in evaluating general performance metrics as recognized in Larson’s original work (Larson 1974). We validate the proposed framework by testing it on both synthetic systems, with and without overlapping regions, and applying it to a real police system in collaboration with the Atlanta Police Department (APD). With the adoption of the “umbrella car” strategy by the APD, where multiple officers collaborate to cover more areas due to understaffing and rising crime rates, our study aims to analyze this operational method and help the police evaluate its effectiveness using our model. The results indicate that deploying more police units does not necessarily enhance policing effectiveness. Instead, overlapping police service regions provide a more flexible and dynamic response to varying call frequencies, significantly improving resource allocation efficiency compared to a rigid, non-overlapping structure.
@@
This paper is organized as follows: We begin with a review of relevant literature on spatial queueing systems and Markov models with similar structures. In Section 3, we introduce a novel queueing framework for evaluating districting plans with overlapping regions. Section 3.1 discusses the problem setup and key assumptions. Section 3.2 presents a generalized hypercube queueing model – the hyperlattice model – for capturing system dynamics, while Section 3.3 introduces an efficient method for estimating its steady-state distribution. Section 3.4 introduces two key performance measures using the hyperlattice model. Lastly, we demonstrate the effectiveness of our framework through experiments on synthetic systems in Section 4 and conclude in Section 5, discussing potential extensions and generalizations.
2 Related works
@@
Spatial queueing systems modeling with overlapping service regions has received limited attention in the literature. However, recent developments in service systems and their increasing complexity have sparked a growing interest in this field, particularly in the pursuit of more efficient dispatch policies. For instance, the concept of overlapping zones is discussed in Bammi (1975), which only considers completely overlapping service regions. This approach may need to address higher crime rates near boundaries. In contrast, our study focuses on partially overlapping service regions, which are more challenging to analyze but better suited for our purposes.
@@
Richard Larson’s paper (Larson 1974) and his co-authored book with Asaf Odoni, “Urban Operations Research” (Larson and Odoni 1981), introduced the hypercube queueing model, a pioneering work in this field. The hypercube queueing model provides a spatial queueing framework for analyzing and evaluating the performance of service systems. It permits servers to overlap into each other’s territories and provide assistance when busy, making it particularly useful for emergency service systems, such as police and ambulance operations (Ansari et al. 2017, de Souza et al. 2015, Takeda et al. 2007, Zhu et al. 2020, 2022, Morabito et al. 2008). However, it assumes identical service from all servers sharing a single queue, which may be oversimplified for real-world service systems and is unable to capture more complex systems. In our context, servers may have limited authorized movement, resulting in separate queues for each server. This is because other servers cannot travel to specific regions to offer assistance. For instance, Beojone et al. (2021) proposed an extension of the hypercube model with fully dedicated servers. While their model allows both dedicated and basic servers to serve the same region under certain random dispatch strategies, it differs significantly from our hyperlattice queueing model in how server responsibilities are handled. In their model, dedicated servers are specifically assigned to handle the most severe calls from any region. In contrast, our hyperlattice queueing model assigns equal responsibilities to all servers, focusing on service regions that can be covered by multiple servers. A similar distinction is found in de Souza et al. (2015), where the model considers queue priorities and assigns different responsibilities to servers, such as Advanced Life Support (ALS) units and Basic Life Support (BLS) units. In their models, overlapping service occurs only across different server types. Moreover, Takeda et al. (2007) considered decentralized but dedicated service areas for ambulances, with each ambulance assigned to a specific station and region. In their hypercube model, overlapping service is achieved by enlarging the service regions with decentralized units. In contrast, our hyperlattice model allows units to serve neighboring overlapping regions without requiring decentralization.
@@
Among existing works extending the classic hypercube queueing model for specific applications, a notable contribution is Beojone et al. (2021), which introduces a variation of the hypercube model. Their extension is designed to represent a dispatch policy in which advanced-equipped servers serve solely life-threatening calls (called dedicated servers). Our model, however, differs as it allows servers to operate within the territories of other servers, presenting a distinct problem framework. Additionally, they grouped servers with similar characteristics in their state space representation, aiming to reduce computational demands. In our method, each server is still represented individually in the state space, managing to do so without significant computational expense. We also assume that each server has its own queue, diverging from most hypercube model variations, which typically use a single queue for all servers. A detailed comparison between various hypercube models and our hyperlattice model is presented in Table 1.
Table 1: Comparison of Hyperlattice Model with Hypercube Models
Model Feature Larson (1974) Beojone et al. (2021) Our Model
Service Region Fully overlapping Fully overlapping Flexible overlapping
Call Priority Uniform Tiered (Urgent/Basic) Uniform
Dispatch Policy Priority-based Priority-based Flexible (state-dependent)
Queueing Structure Single loss queue Single general queue Multiple general queue
Tractability High Moderate High (via truncation)
@@
Recently, Tsitsiklis and Xu (2017) introduced a flexible queue architecture for a multi-server model, comprising $n$ flexible servers and $n$ queues connected through a bipartite graph, with each server having its own associated queue. The paper primarily focuses on the model’s theoretical aspects and highlights the key benefits of flexibility and resource pooling in such queueing networks. They specifically concentrate on the scaling regime where the system size $n$ tends to be infinite while the overall traffic intensity remains constant. However, their model cannot be directly extended to our scenario, as it does not account for the geographical relationships among servers.
The queueing literature extensively examines load-balancing models involving multiple dispatchers and flexibility constraints, often under the framework of flexible queues or load-balancing systems. A representative example is the “Join the Idle Queue” policy and related dispatching strategies designed to optimize system performance under varying workload conditions (see, e.g., Cruise et al. (2020), Weng et al. (2020)). These models typically focus on analyzing optimal or near-optimal dispatching policies under simplified assumptions. However, our contribution is primarily modeling-oriented: We develop a generalized hypercube queueing framework that captures the spatial and structural complexity of real-world service systems. Unlike much of the load-balancing literature, which often assumes homogeneous servers, centralized coordination, or simplified system dynamics, our model explicitly incorporates server-level queues, spatially dependent service areas, and flexible, state-dependent dispatch policies. Instead of focusing on specific policies, we develop a general and extensible framework that approximates system performance across a broad class of state-dependent dispatch policies, while maintaining computational tractability.